OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

OpenStudy (anmolsingh):

3x-7=5x+19 now get all the "x" on one side 5x-3x=-7-19 2x=-26 x=-13

OpenStudy (anmolsingh):

now you can check the answer by substituting the value of -13 in the equation and see if both the sides are equal

OpenStudy (anonymous):